The mode cleaner is a little misaligned in pitch, and is very misaligned in yaw. The lowest order mode that is flashing is TEM11.
I had a look-see at the SUS sensors, to see if there were any big jumps. There were moderately sized jumps on all 3 mode cleaner optics.

The MC's lockloss was at ~8:22am this morning, and went along with a giganto kick to the optics. Steve tells me that Den might have been kicking up optics while doing computer things this morning, before Steve reminded him to shut off the watchdogs. However, Steve was also taking phots/measuring things near MC Refl, so maybe he's not totally absolved of blame. But this really looks like the optics settled to different places after big kicks.
I'm going to try to align the MC mirrors to get back to the sensor numbers from early this morning before chaos began.
Reminder / Moral: Everything cannot be considered to be "working fine" if the MC isn't locking. See if you can figure out why, and especially if it's something that you screwed up, either fix it, or better yet, ask for help and learn how to fix what you broke. |